Hello,
Can someone please help me find the x-intercept in 3x+y=4?

Thank you

Answer by ptaylor(1894)

Can someone please help me find the x-intercept in 3x+y=4?

This is the equation of a straight line
The x intercept is found by letting y=0
3x+0=4 divide both sides by 3
x=4/3----------------x-intercept


Hope this helps----ptaylor